The order reads: Dopamine 10 mcg/kg/min. The bag is labeled 800mg in 250mL of NS. The patient weighs 242 pounds.   How many mL/hour will be infused?
Answer #1

Ordered Dopamine= 10 mcg/kg/min

weight of the patient = 242 pound= 242/2.2=110 kg

hence as per the weight of the patient required dose will be= 110 X 10 mcg/min

= 1100 mcg/min

= 1.1 mg/min as (1 mg= 1000 mcg)

Labeled 800mg in 250mL

Means 800/250

= 3.2 mg/ml

3.2 mg in 1 ml

1 mg in = 1/3.2 ml

hence 1.1 mg will be in = 1.1/3.2 ml=0.3437 ml

Hence the required dose will be = 0.3437 ml/min

Required dose in ml/hr = 20.625 ml/hour